# Recursive construction and enumeration of self-orthogonal and self-dual codes over finite commutative chain rings of even characteristic - I

### v1

@@ 2 Some preliminaries
      $(2^{m})^{\sum\limits_{i=\ell}^{\gamma_{\ell}+\ell}\lambda_{i}\Lambda_{i-\ell+1}+\Lambda_{\gamma_{\ell}+1}(n-\Lambda_{\gamma_{\ell}+\ell})-\Lambda_{s-2\mathrm{f}_{\ell}+2}-\Lambda_{\gamma_{\ell}+1-\kappa_{1}}-\frac{\Lambda_{\gamma_{\ell}+1}(\Lambda_{\gamma_{\ell}+1}-1)}{2}+1}.$
      Next, for a given choice of the matrix $V^{(\ell-2)}]_{\gamma_{\ell}+1}$ satisfying (3.24)-(3.26), we can write
@@
      $(2^{m})^{\sum\limits_{i=\ell+1}^{\gamma_{\ell}+\ell}\lambda_{i}\Lambda_{i-\ell}+\Lambda_{\gamma_{\ell}+1}(n-\Lambda_{\gamma_{\ell}+\ell})-\Lambda_{s-2\mathrm{f}_{\ell}+1}-\frac{\Lambda_{\gamma_{\ell}+1}(\Lambda_{\gamma_{\ell}+1}-1)}{2}}$
@@
      $(2^{m})^{(\Lambda_{\gamma_{\ell}+\ell-1}-\Lambda_{\gamma_{\ell}+1})(n-\Lambda_{\gamma_{\ell}+\ell}-\Lambda_{\gamma_{\ell}+1})}{\lambda_{\gamma_{\ell}+\ell}+n-\Lambda_{\gamma_{\ell}+\ell}-\Lambda_{\gamma_{\ell}+1}\brack\lambda_{\gamma_{\ell}+\ell}}_{2^{m}}$
@@
The following proposition lifts a self-orthogonal code of type $\{\Lambda_{s-\kappa_{1}+1},\lambda_{s-\kappa_{1}+2},\ldots,\lambda_{s+\kappa_{1}}\}$ and length $n$ over $\mathscr{R}_{\kappa-1,m}$ satisfying property $(\mathfrak{P})$ to a self-orthogonal code of type $\{\Lambda_{s-\kappa_{1}},\lambda_{s-\kappa_{1}+1},\ldots,\lambda_{s+\kappa_{1}+1}\}$ and the same length $n$ over $\mathscr{R}_{\kappa+1,m}$ satisfying property $(\mathfrak{P})$ when $e$ is even and $2\kappa\leq e.$ It also enumerates the number of possible ways to perform this lifting.
@@
Let $e$ be even, $2\kappa\leq e,$ and let $\ell=\kappa+1.$ Let $\mathscr{D}_{\kappa-1}$ be a self-orthogonal code of type $\{\Lambda_{s-\kappa_{1}+1},\lambda_{s-\kappa_{1}+2},\ldots,\lambda_{s+\kappa_{1}}\}$ and length $n$ over $\mathscr{R}_{\kappa-1,m}$ satisfying property $(\mathfrak{P}).$ Let $\mathcal{C}^{(s-\kappa)}\subseteq\mathcal{C}^{(s-\kappa+1)}\subseteq\mathcal{C}^{(s-\kappa_{1})}$ be a chain of linear subcodes of $Tor_{1}(\mathscr{D}_{\kappa-1}),$ such that $\dim\mathcal{C}^{(s-\kappa)}=\Lambda_{s-\kappa},$ $\dim\mathcal{C}^{(s-\kappa+1)}=\Lambda_{s-\kappa+1}$, $\dim\mathcal{C}^{(s-\kappa_{1})}=\Lambda_{s-\kappa_{1}}$ and the code $\mathcal{C}^{(s-\kappa_{1})}$ is doubly even, (such a chain $\mathcal{C}^{(s-\kappa)}\subseteq\mathcal{C}^{(s-\kappa+1)}\subseteq\mathcal{C}^{(s-\kappa_{1})}$ of codes exists for each choice of $\mathscr{D}_{\kappa-1},$ by Lemma 3.2). The following hold.
@@
    There exists a self-orthogonal code $\mathscr{D}_{\kappa+1}$ of type $\{\Lambda_{s-\kappa_{1}},\lambda_{s-\kappa_{1}+1},\ldots,\lambda_{s+\kappa_{1}+1}\}$ and length $n$ over $\mathscr{R}_{\kappa+1,m}$ satisfying property $(\mathfrak{P})$ with $Tor_{1}(\mathscr{D}_{\kappa+1})=\mathcal{C}^{(s-\kappa_{1})}$ and $Tor_{j+1}(\mathscr{D}_{\kappa+1})=Tor_{j}(\mathscr{D}_{\kappa-1})$ for $1\leq j\leq\kappa-1,$ where such a code exists:
    • unconditionally, if $\textbf{1}\notin\mathcal{C}^{(s-\kappa)};$
    • if and only if either $n\equiv 0\pmod{8}$ or $n\equiv 4\pmod{8}$ with $m$ even when $\textbf{1}\in\mathcal{C}^{(s-\kappa)}.$
@@
    Moreover, each such choice of the codes $\mathscr{D}_{\kappa-1},$ $\mathcal{C}^{(s-\kappa_{1})},$ $\mathcal{C}^{(s-\kappa+1)}$ and $\mathcal{C}^{(s-\kappa)}$ gives rise to precisely
    $2^{\epsilon}(2^{m})^{\sum\limits_{i=\kappa+1}^{s+\kappa_{1}+1}\lambda_{i}\Lambda_{i-\kappa}+\sum\limits_{j=\kappa+2}^{s+\kappa_{1}+1}\lambda_{j}\Lambda_{j-\kappa-1}+Y_{\kappa+1}}{\lambda_{s+\kappa_{1}+1}+n-\Lambda_{s+\kappa_{1}+1}-\Lambda_{s-\kappa_{1}}\brack\lambda_{s+\kappa_{1}+1}}_{2^{m}}$
    distinct self-orthogonal codes $\mathscr{D}_{\kappa+1}$ of type $\{\Lambda_{s-\kappa_{1}},\lambda_{s-\kappa_{1}+1},\ldots,\lambda_{s+\kappa_{1}+1}\}$ and length $n$ over $\mathscr{R}_{\kappa+1,m}$ satisfying property $(\mathfrak{P})$ with $Tor_{1}(\mathscr{D}_{\kappa+1})=\mathcal{C}^{(s-\kappa_{1})}$ and $Tor_{j+1}(\mathscr{D}_{\kappa+1})=Tor_{j}(\mathscr{D}_{\kappa-1})$ for $1\leq j\leq\kappa-1,$ where $\epsilon=1$ if $\textbf{1}\in\mathcal{C}^{(s-\kappa)},$ while $\epsilon=0$ otherwise, and $Y_{\kappa+1}=(\Lambda_{s+\kappa_{1}}+\Lambda_{s-\kappa_{1}})(n-\Lambda_{s+\kappa_{1}+1}-\Lambda_{s-\kappa_{1}})+\Lambda_{s-\kappa_{1}}^{2}+\Lambda_{s-\kappa_{1}}-\Lambda_{s-\kappa}-2\Lambda_{s-\kappa+1}+\omega_{\kappa+1}$ with $\omega_{\kappa+1}=1$ if $\textbf{1}\in\mathcal{C}^{(s-\kappa+1)},$ while $\omega_{\kappa+1}=0$ otherwise.
@@
To prove the result, we assume, without any loss of generality, that the code $\mathscr{D}_{\kappa-1}$ has a generator matrix
@@
$\mathtt{W}^{(0)}]_{s-\kappa_{1}+1}=\begin{bmatrix}\mathtt{W}_{1}^{(0)}\\ \mathtt{W}_{2}^{(0)}\\ \vdots\\ \mathtt{W}_{s-\kappa_{1}+1}^{(0)}\end{bmatrix}=\begin{bmatrix}{\mathtt{I}}_{\lambda_{1}}&{\mathtt{B}}_{1,1}^{(0)}&\cdots&{\mathtt{B}}_{1,s-\kappa_{1}}^{(0)}&\cdots&{\mathtt{B}}_{1,e-1}^{(0)}&{\mathtt{B}}_{1,e}^{(0)}\\ \mathbf{0}&{\mathtt{I}}_{\lambda_{2}}&\cdots&{\mathtt{B}}_{2,s-\kappa_{1}}^{(0)}&\cdots&{\mathtt{B}}_{2,e-1}^{(0)}&{\mathtt{B}}_{2,e}^{(0)}\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ \mathbf{0}&\mathbf{0}&\cdots&{\mathtt{I}}_{\lambda_{s-\kappa_{1}+1}}&\cdots&{\mathtt{B}}_{s-\kappa_{1}+1,e-1}^{(0)}&{\mathtt{B}}_{s-\kappa_{1}+1,e}^{(0)}\end{bmatrix}$
with ${\mathtt{I}}_{\lambda_{i}}$ as the $\lambda_{i}\times\lambda_{i}$ identity matrix over ${\mathcal{T}}_{m},$ ${\mathtt{B}}_{i,j}^{(0)}\in\mathcal{M}_{\lambda_{i}\times\lambda_{j+1}}({\mathcal{T}}_{m})$ for $1\leq i\leq s-\kappa_{1}+1$ and $i\leq j\leq e,$ $\mathtt{V}^{(a)}]_{s-\kappa_{1}+1}\in\mathcal{M}_{\Lambda_{s-\kappa_{1}+1}\times n}(\mathcal{T}_{m})$ for $1\leq a\leq\kappa-2,$ and the matrix $\mathtt{W}_{s-\kappa_{1}-1+g}^{(\kappa-1)}\in\mathcal{M}_{\lambda_{s-\kappa_{1}-1+g}\times n}(\mathscr{R}_{\kappa-1,m})$ to be considered modulo $u^{\kappa+1-g}$ for $3\leq g\leq\kappa.$ Note that the torsion code $Tor_{1}(\mathscr{D}_{\kappa-1})$ is a $\Lambda_{s-\kappa_{1}+1}$-dimensional code over $\mathcal{T}_{m}$ with a generator matrix $\mathtt{W}^{(0)}]_{s-\kappa_{1}+1}.$ Here, we also assume, without any loss of generality, that the code $\mathcal{C}^{(s-\kappa_{1})}$ has a generator matrix $\mathtt{W}^{(0)}]_{s-\kappa_{1}},$ the code $\mathcal{C}^{(s-\kappa+1)}$ has a generator matrix $\mathtt{W}^{(0)}]_{s-\kappa+1}$ and the code $\mathcal{C}^{(s-\kappa)}$ has a generator matrix $\mathtt{W}^{(0)}]_{s-\kappa}.$ Moreover, by Remark 3.1, we assume that the matrix $(\mathtt{B}^{(0)})_{s-\kappa_{1},s+\kappa_{1}+1}$ has full row-rank.
Since $\mathscr{D}_{\kappa-1}$ is a self-orthogonal code over $\mathscr{R}_{\kappa-1,m}$ satisfying property $(\mathfrak{P}),$ we have
@@
$\mathtt{G}_{\kappa+1}=\begin{bmatrix}\mathtt{W}_{1}^{(\kappa+1)}\\ \mathtt{W}_{2}^{(\kappa+1)}\\ \vdots\\ \mathtt{W}_{s-\kappa_{1}}^{(\kappa+1)}\\ \vskip 1.42262ptu\mathtt{W}_{s-\kappa_{1}+1}^{(\kappa+1)}\\ \vdots\\ u^{\kappa}\mathtt{W}_{s+\kappa_{1}+1}^{(\kappa+1)}\end{bmatrix}=\begin{bmatrix}\mathtt{W}_{1}^{(\kappa-1)}+u^{\kappa-1}\mathtt{V}_{1}^{(\kappa-1)}+u^{\kappa}\mathtt{V}_{1}^{(\kappa)}\\ \mathtt{W}_{2}^{(\kappa-1)}+u^{\kappa-1}\mathtt{V}_{2}^{(\kappa-1)}+u^{\kappa}\mathtt{V}_{2}^{(\kappa)}\\ \vdots\\ \mathtt{W}_{s-\kappa_{1}}^{(\kappa-1)}+u^{\kappa-1}\mathtt{V}_{s-\kappa_{1}}^{(\kappa-1)}+u^{\kappa}\mathtt{V}_{s-\kappa_{1}}^{(\kappa)}\\ \vskip 1.42262ptu\mathtt{W}_{s-\kappa_{1}+1}^{(\kappa+1)}\\ \vdots\\ u^{\kappa}\mathtt{W}_{s+\kappa_{1}+1}^{(\kappa+1)}\end{bmatrix}$,
@@
with $\mathtt{B}_{i,j}^{(\mu)}\in\mathcal{M}_{\lambda_{i}\times\lambda_{j+1}}(\mathcal{T}_{m})$ for $1\leq i\leq s-\kappa_{1}$ and $i+\mu\leq j\leq e,$ $\mathtt{B}_{s-\kappa_{1}-1+g,h}^{(\kappa+1-g)}\in\mathcal{M}_{\lambda_{s-\kappa_{1}-1+g}\times\lambda_{h+1}}(\mathcal{T}_{m})$ for $s+\kappa_{1}+1\leq h\leq e,$ and $\mathtt{B}_{s+\kappa_{1}+1,a}^{(0)}\in\mathcal{M}_{\lambda_{s+\kappa_{1}+1}\times\lambda_{a+1}}(\mathcal{T}_{m})$ for $s+\kappa_{1}+1\leq a\leq e.$ Let $\mathscr{D}_{\kappa+1}$ be a linear code of length $n$ over $\mathscr{R}_{\kappa+1,m}$ with a generator matrix $\mathtt{G}_{\kappa+1}.$ Note that the code $\mathscr{D}_{\kappa+1}$ is of type $\{\Lambda_{s-\kappa_{1}},\lambda_{s-\kappa_{1}+1},\ldots,\lambda_{s+\kappa_{1}+1}\}$ over $\mathscr{R}_{\kappa+1,m}$ satisfying $Tor_{1}(\mathscr{D}_{\kappa+1})=\mathcal{C}^{(s-\kappa_{1})}$ and $Tor_{i+1}(\mathscr{D}_{\kappa+1})=Tor_{i}(\mathscr{D}_{\kappa-1})$ for $1\leq i\leq\kappa-1.$ Moreover, by Lemma 2.1, we see that the code $\mathscr{D}_{\kappa+1}$ is a self-orthogonal code over $\mathscr{R}_{\kappa+1,m}$ satisfying property $(\mathfrak{P})$ if and only if there exist matrices $\mathtt{V}^{(\kappa-1)}]_{s-\kappa_{1}}, [$$\mathtt{V}^{(\kappa)}]_{s-\kappa_{1}}, [$$\left[\mathbf{0}~\cdots~\mathbf{0}~\mathtt{B}_{s-\kappa_{1}-1+g,s+\kappa_{1}+1}^{(\kappa+1-g)}~\cdots~\mathtt{B}_{s-\kappa_{1}-1+g,e}^{(\kappa+1-g)}\right]$ for $2\leq g\leq\kappa$ and $\mathtt{W}_{s+\kappa_{1}+1}^{(\kappa+1)}$ satisfying the following system of matrix equations:
@@
We will now establish the existence of the matrices $\mathtt{V}^{(\kappa-1)}]_{s-\kappa_{1}}, [$$\mathtt{V}^{(\kappa)}]_{s-\kappa_{1}}, [$$\Big[\mathbf{0}~\cdots~\mathbf{0}~\mathtt{B}_{s-\kappa_{1}-1+g,s+\kappa_{1}+1}^{(\kappa+1-g)}~\cdots~\\ \mathtt{B}_{s-\kappa_{1}-1+g,e}^{(\kappa+1-g)}\Big]$ for $2\leq g\leq\kappa$ and $\mathtt{W}_{s+\kappa_{1}+1}^{(\kappa+1)}$ satisfying the system of matrix equations (3.39)-(3.44) and count their choices. To do this, we first recall that the matrix $(\mathtt{B}^{(0)})_{s-\kappa_{1},s+\kappa_{1}+1}$ has full row-rank over $\mathcal{T}_{m}.$ Working as in Proposition 3.3, we see that there exists a matrix $\mathtt{V}^{(\kappa-1)}]_{s-\kappa_{1}}$ satisfying the following system of matrix equations:

### v2

@@ 2.6 Notations and preliminary setup
      $\vskip-2.84526pt\displaystyle(2^{m})^{\sum\limits_{i=\ell}^{\gamma_{\ell}+\ell}\lambda_{i}\Lambda_{i-\ell+1}+\Lambda_{\gamma_{\ell}+1}(n-\Lambda_{\gamma_{\ell}+\ell})-\Lambda_{s-2\mathrm{f}_{\ell}+2}-\Lambda_{\gamma_{\ell}+1-\kappa_{1}}-\frac{\Lambda_{\gamma_{\ell}+1}(\Lambda_{\gamma_{\ell}+1}-1)}{2}+1}.$
      Next, for a given choice of the matrix $V^{(\ell-2)}]_{\gamma_{\ell}+1}$ satisfying (3.24)–(3.26), we can write
@@
      $\vskip-2.84526pt\displaystyle(2^{m})^{\sum\limits_{i=\ell+1}^{\gamma_{\ell}+\ell}\lambda_{i}\Lambda_{i-\ell}+\Lambda_{\gamma_{\ell}+1}(n-\Lambda_{\gamma_{\ell}+\ell})-\Lambda_{s-2\mathrm{f}_{\ell}+1}-\frac{\Lambda_{\gamma_{\ell}+1}(\Lambda_{\gamma_{\ell}+1}-1)}{2}}$
@@
      $\vskip-2.84526pt(2^{m})^{(\Lambda_{\gamma_{\ell}+\ell-1}-\Lambda_{\gamma_{\ell}+1})(n-\Lambda_{\gamma_{\ell}+\ell}-\Lambda_{\gamma_{\ell}+1})}{\lambda_{\gamma_{\ell}+\ell}+n-\Lambda_{\gamma_{\ell}+\ell}-\Lambda_{\gamma_{\ell}+1}\brack\lambda_{\gamma_{\ell}+\ell}}_{2^{m}}$
@@
The following proposition lifts a self-orthogonal code of type $\{\Lambda_{s-\kappa_{1}+1},\lambda_{s-\kappa_{1}+2},\ldots,\lambda_{s+\kappa_{1}+\theta_{e}}\}$ and length $n$ over $\mathscr{R}_{\kappa-1+\theta_{e},m}$ satisfying property $(\mathfrak{P})$ to a self-orthogonal code of type $\{\Lambda_{s-\kappa_{1}},\lambda_{s-\kappa_{1}+1},\ldots,\lambda_{s+\kappa_{1}+1+\theta_{e}}\}$ and the same length $n$ over $\mathscr{R}_{\kappa+1+\theta_{e},m}$ that also satisfies property $(\mathfrak{P})$ when $2\kappa\leq e.$ It also enumerates the number of all possible distinct ways to perform this lifting.
@@
Let $2\kappa\leq e,$ and let $\ell=\kappa+1+\theta_{e}.$ Let $\mathscr{D}_{\kappa-1+\theta_{e}}$ be a self-orthogonal code of type $\{\Lambda_{s-\kappa_{1}+1},\lambda_{s-\kappa_{1}+2},\\ \ldots,\lambda_{s+\kappa_{1}+\theta_{e}}\}$ and length $n$ over $\mathscr{R}_{\kappa-1+\theta_{e},m}$ satisfying property $(\mathfrak{P}).$ Let $\mathcal{C}^{(s-\kappa+\theta_{e})}\subseteq\mathcal{C}^{(s-\kappa+1)}\subseteq\mathcal{C}^{(s-\kappa_{1})}$ be a chain of linear subcodes of $Tor_{1}(\mathscr{D}_{\kappa-1+\theta_{e}}),$ such that $\dim\mathcal{C}^{(s-\kappa+\theta_{e})}=\Lambda_{s-\kappa+\theta_{e}},$ $\dim\mathcal{C}^{(s-\kappa+1)}=\Lambda_{s-\kappa+1}$, $\dim\mathcal{C}^{(s-\kappa_{1})}=\Lambda_{s-\kappa_{1}}$ and the code $\mathcal{C}^{(s-\kappa_{1})}$ is doubly even, (such a chain $\mathcal{C}^{(s-\kappa+\theta_{e})}\subseteq\mathcal{C}^{(s-\kappa+1)}\subseteq\mathcal{C}^{(s-\kappa_{1})}$ of codes exists for each choice of $\mathscr{D}_{\kappa-1+\theta_{e}},$ by Lemma 3.2). The following hold.
@@
    There exists a self-orthogonal code $\mathscr{D}_{\kappa+1+\theta_{e}}$ of type $\{\Lambda_{s-\kappa_{1}},\lambda_{s-\kappa_{1}+1},\ldots,\lambda_{s+\kappa_{1}+1+\theta_{e}}\}$ and length $n$ over $\mathscr{R}_{\kappa+1+\theta_{e},m}$ satisfying property $(\mathfrak{P})$ with $Tor_{1}(\mathscr{D}_{\kappa+1+\theta_{e}})=\mathcal{C}^{(s-\kappa_{1})}$ and $Tor_{j+1}(\mathscr{D}_{\kappa+1+\theta_{e}})=Tor_{j}(\mathscr{D}_{\kappa-1+\theta_{e}})$ for $1\leq j\leq\kappa-1+\theta_{e},$ where such a code exists:
    • unconditionally, if $\textbf{1}\notin\mathcal{C}^{(s-\kappa+\theta_{e})};$
    • if and only if either $n\equiv 0\pmod{8}$ or $n\equiv 4\pmod{8}$ with $m$ even when $\textbf{1}\in\mathcal{C}^{(s-\kappa+\theta_{e})}.$
@@
    Moreover, each such choice of the codes $\mathscr{D}_{\kappa-1+\theta_{e}},$ $\mathcal{C}^{(s-\kappa_{1})},$ $\mathcal{C}^{(s-\kappa+1)}$ and $\mathcal{C}^{(s-\kappa+\theta_{e})}$ gives rise to precisely
    $\vskip-2.84526pt\hskip-11.38109pt2^{\epsilon}(2^{m})^{\sum\limits_{i=\kappa+1+\theta_{e}}^{s+\kappa_{1}+1+\theta_{e}}\lambda_{i}\Lambda_{i-\kappa-\theta_{e}}+\sum\limits_{j=\kappa+2+\theta_{e}}^{s+\kappa_{1}+1+\theta_{e}}\lambda_{j}\Lambda_{j-\kappa-1-\theta_{e}}+\Lambda_{s-\kappa_{1}}^{2}+\Lambda_{s-\kappa_{1}}+Y_{\kappa+1+\theta_{e}}}{\lambda_{s+\kappa_{1}+1+\theta_{e}}+n-\Lambda_{s+\kappa_{1}+1+\theta_{e}}-\Lambda_{s-\kappa_{1}}\brack\lambda_{s+\kappa_{1}+1+\theta_{e}}}_{2^{m}}$
    distinct self-orthogonal codes $\mathscr{D}_{\kappa+1+\theta_{e}}$ of type $\{\Lambda_{s-\kappa_{1}},\lambda_{s-\kappa_{1}+1},\ldots,\lambda_{s+\kappa_{1}+1+\theta_{e}}\}$ and length $n$ over $\mathscr{R}_{\kappa+1+\theta_{e},m}$ satisfying property $(\mathfrak{P})$ with $Tor_{1}(\mathscr{D}_{\kappa+1+\theta_{e}})=\mathcal{C}^{(s-\kappa_{1})}$ and $Tor_{j+1}(\mathscr{D}_{\kappa+1+\theta_{e}})=Tor_{j}(\mathscr{D}_{\kappa-1+\theta_{e}})$ for $1\leq j\leq\kappa-1+\theta_{e}.$ Here $\epsilon=1$ if $\textbf{1}\in\mathcal{C}^{(s-\kappa+\theta_{e})},$ while $\epsilon=0$ otherwise. The number $Y_{\kappa+1+\theta_{e}}$ is given by Y_κ+1+θ_e={ (Λ_s+κ_1+1+Λ_s-κ_1)(n-Λ_s+κ_1+2-Λ_s-κ_1)-Λ_s-κ+1-Λ_s-κif e is odd;(Λ_s+κ_1+Λ_s-κ_1)(n-Λ_s+κ_1+1-Λ_s-κ_1)-Λ_s-κ -2Λ_s-κ+1 +ω_κ+1if e is even,\@unrecurse where, in the case that $e$ is even, the number $\omega_{\kappa+1}$ is given by $\omega_{\kappa+1}=1$ if $\textbf{1}\in\mathcal{C}^{(s-\kappa+1)},$ while $\omega_{\kappa+1}=0$ otherwise.
@@
To prove the result, we will consider the following two cases separately: (I) $e$ is even , and (II) $e$ is odd.
  • (I)
    Let $e$ be even. In this case, we assume, without any loss of generality, that the code $\mathscr{D}_{\kappa-1}$ has a generator matrix
@@
    $\vskip-2.84526pt[\mathtt{W}^{(0)}]_{s-\kappa_{1}+1}=\begin{bmatrix}\mathtt{W}_{1}^{(0)}\\ \mathtt{W}_{2}^{(0)}\\ \vdots\\ \mathtt{W}_{s-\kappa_{1}+1}^{(0)}\end{bmatrix}=\begin{bmatrix}{\mathtt{I}}_{\lambda_{1}}&{\mathtt{B}}_{1,1}^{(0)}&\cdots&{\mathtt{B}}_{1,s-\kappa_{1}}^{(0)}&\cdots&{\mathtt{B}}_{1,e-1}^{(0)}&{\mathtt{B}}_{1,e}^{(0)}\\ \mathbf{0}&{\mathtt{I}}_{\lambda_{2}}&\cdots&{\mathtt{B}}_{2,s-\kappa_{1}}^{(0)}&\cdots&{\mathtt{B}}_{2,e-1}^{(0)}&{\mathtt{B}}_{2,e}^{(0)}\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ \mathbf{0}&\mathbf{0}&\cdots&{\mathtt{I}}_{\lambda_{s-\kappa_{1}+1}}&\cdots&{\mathtt{B}}_{s-\kappa_{1}+1,e-1}^{(0)}&{\mathtt{B}}_{s-\kappa_{1}+1,e}^{(0)}\end{bmatrix}$
    with ${\mathtt{I}}_{\lambda_{i}}$ as the $\lambda_{i}\times\lambda_{i}$ identity matrix over ${\mathcal{T}}_{m},$ ${\mathtt{B}}_{i,j}^{(0)}\in\mathcal{M}_{\lambda_{i}\times\lambda_{j+1}}({\mathcal{T}}_{m})$ for $1\leq i\leq s-\kappa_{1}+1$ and $i\leq j\leq e,$ $\mathtt{V}^{(a)}]_{s-\kappa_{1}+1}\in\mathcal{M}_{\Lambda_{s-\kappa_{1}+1}\times n}(\mathcal{T}_{m})$ for $1\leq a\leq\kappa-2,$ and the matrix $\mathtt{W}_{s-\kappa_{1}-1+g}^{(\kappa-1)}\in\mathcal{M}_{\lambda_{s-\kappa_{1}-1+g}\times n}(\mathscr{R}_{\kappa-1,m})$ to be considered modulo $u^{\kappa+1-g}$ for $3\leq g\leq\kappa.$ Note that the torsion code $Tor_{1}(\mathscr{D}_{\kappa-1})$ is a $\Lambda_{s-\kappa_{1}+1}$-dimensional code over $\mathcal{T}_{m}$ with a generator matrix $\mathtt{W}^{(0)}]_{s-\kappa_{1}+1}.$ Here, we also assume, without any loss of generality, that the code $\mathcal{C}^{(s-\kappa_{1})}$ has a generator matrix $\mathtt{W}^{(0)}]_{s-\kappa_{1}},$ the code $\mathcal{C}^{(s-\kappa+1)}$ has a generator matrix $\mathtt{W}^{(0)}]_{s-\kappa+1}$ and the code $\mathcal{C}^{(s-\kappa)}$ has a generator matrix $\mathtt{W}^{(0)}]_{s-\kappa}.$ Moreover, by Remark 3.1, we assume that the matrix $(\mathtt{B}^{(0)})_{s-\kappa_{1},s+\kappa_{1}+1}$ has full row-rank. Since $\mathscr{D}_{\kappa-1}$ is a self-orthogonal code over $\mathscr{R}_{\kappa-1,m}$ satisfying property $(\mathfrak{P}),$ we have
@@
    $\vskip-2.84526pt\mathtt{G}_{\kappa+1}=\begin{bmatrix}\mathtt{W}_{1}^{(\kappa+1)}\\ \mathtt{W}_{2}^{(\kappa+1)}\\ \vdots\\ \mathtt{W}_{s-\kappa_{1}}^{(\kappa+1)}\\ \vskip 1.42262ptu\mathtt{W}_{s-\kappa_{1}+1}^{(\kappa+1)}\\ \vdots\\ u^{\kappa}\mathtt{W}_{s+\kappa_{1}+1}^{(\kappa+1)}\end{bmatrix}=\begin{bmatrix}\mathtt{W}_{1}^{(\kappa-1)}+u^{\kappa-1}\mathtt{V}_{1}^{(\kappa-1)}+u^{\kappa}\mathtt{V}_{1}^{(\kappa)}\\ \mathtt{W}_{2}^{(\kappa-1)}+u^{\kappa-1}\mathtt{V}_{2}^{(\kappa-1)}+u^{\kappa}\mathtt{V}_{2}^{(\kappa)}\\ \vdots\\ \mathtt{W}_{s-\kappa_{1}}^{(\kappa-1)}+u^{\kappa-1}\mathtt{V}_{s-\kappa_{1}}^{(\kappa-1)}+u^{\kappa}\mathtt{V}_{s-\kappa_{1}}^{(\kappa)}\\ \vskip 1.42262ptu\mathtt{W}_{s-\kappa_{1}+1}^{(\kappa+1)}\\ \vdots\\ u^{\kappa}\mathtt{W}_{s+\kappa_{1}+1}^{(\kappa+1)}\end{bmatrix}$,
@@
    with $\mathtt{B}_{i,j}^{(\mu)}\in\mathcal{M}_{\lambda_{i}\times\lambda_{j+1}}(\mathcal{T}_{m})$ for $1\leq i\leq s-\kappa_{1}$ and $i+\mu\leq j\leq e,$ $\mathtt{B}_{s-\kappa_{1}-1+g,h}^{(\kappa+1-g)}\in\mathcal{M}_{\lambda_{s-\kappa_{1}-1+g}\times\lambda_{h+1}}(\mathcal{T}_{m})$ for $s+\kappa_{1}+1\leq h\leq e,$ and $\mathtt{B}_{s+\kappa_{1}+1,a}^{(0)}\in\mathcal{M}_{\lambda_{s+\kappa_{1}+1}\times\lambda_{a+1}}(\mathcal{T}_{m})$ for $s+\kappa_{1}+1\leq a\leq e.$ Let $\mathscr{D}_{\kappa+1}$ be a linear code of length $n$ over $\mathscr{R}_{\kappa+1,m}$ with a generator matrix $\mathtt{G}_{\kappa+1}.$ Note that the code $\mathscr{D}_{\kappa+1}$ is of type $\{\Lambda_{s-\kappa_{1}},\lambda_{s-\kappa_{1}+1},\ldots,\lambda_{s+\kappa_{1}+1}\}$ over $\mathscr{R}_{\kappa+1,m}$ satisfying $Tor_{1}(\mathscr{D}_{\kappa+1})=\mathcal{C}^{(s-\kappa_{1})}$ and $Tor_{i+1}(\mathscr{D}_{\kappa+1})=Tor_{i}(\mathscr{D}_{\kappa-1})$ for $1\leq i\leq\kappa-1.$ Moreover, by Lemma 2.1, one can easily observe that the code $\mathscr{D}_{\kappa+1}$ is a self-orthogonal code over $\mathscr{R}_{\kappa+1,m}$ satisfying property $(\mathfrak{P})$ if and only if there exist matrices $\mathtt{V}^{(\kappa-1)}]_{s-\kappa_{1}}, [$$\mathtt{V}^{(\kappa)}]_{s-\kappa_{1}}, [$$\mathbf{0}~\mathbf{0}~\cdots~\mathbf{0}~\mathtt{B}_{s-\kappa_{1}-1+g,s+\kappa_{1}+1}^{(\kappa+1-g)}~\cdots~\mathtt{B}_{s-\kappa_{1}-1+g,e}^{(\kappa+1-g)}] [$$\left[\mathbf{0}~\mathbf{0}~\cdots~\mathbf{0}~\mathtt{B}_{s-\kappa_{1}-1+g,s+\kappa_{1}+1}^{(\kappa+1-g)}~\cdots~\mathtt{B}_{s-\kappa_{1}-1+g,e}^{(\kappa+1-g)}\right]$ for $2\leq g\leq\kappa$ and $\mathtt{W}_{s+\kappa_{1}+1}^{(\kappa+1)}$ satisfying the following system of matrix equations:
@@
    We will now establish the existence of the matrices $\mathtt{V}^{(\kappa-1)}]_{s-\kappa_{1}}, [$$\mathtt{V}^{(\kappa)}]_{s-\kappa_{1}}, [$$\Big[\mathbf{0}~\cdots~\mathbf{0}~\mathtt{B}_{s-\kappa_{1}-1+g,s+\kappa_{1}+1}^{(\kappa+1-g)}~\cdots~\\ \mathtt{B}_{s-\kappa_{1}-1+g,e}^{(\kappa+1-g)}\Big]$ for $2\leq g\leq\kappa$ and $\mathtt{W}_{s+\kappa_{1}+1}^{(\kappa+1)}$ satisfying the system of matrix equations (3.39)–(3.44) and count their choices. To do this, we first recall that the matrix $(\mathtt{B}^{(0)})_{s-\kappa_{1},s+\kappa_{1}+1}$ has full row-rank over $\mathcal{T}_{m}.$ Working as in Proposition 3.3, we see that there exists a matrix $\mathtt{V}^{(\kappa-1)}]_{s-\kappa_{1}}$ satisfying the following system of matrix equations: